OpenStudy (anonymous):

decide whether the lines are parallell, perpendicular, or neither 5x+3y=10 3x+5y=4

OpenStudy (aroub):

3y=-5x+10 y=-5/3x+10/3 5y=-3x+4 y=-3/5x+4/5 so it's neither

OpenStudy (anonymous):

thanks